OpenStudy (anonymous):

Solve the system of equations. x + 3y = -1 2x + 2y = 6 A.(-4,1) B.(2,-1) C.(4,-1) D.(5,-2)

OpenStudy (anonymous):

Multiply the first equation by 2 so that x+3y=-1 becomes 2x+6y=-2. Next, you subtract them to get 0x+4y=-8, or 4y=-8. Divide both sides by 4 to get y=-2. Next, you simply substitute -2 for y in x+3y=-1 to get x.

OpenStudy (anonymous):

mcfizzy are you in flvs

OpenStudy (anonymous):

the answer is d

OpenStudy (anonymous):

oh sorry i just came back thanks you guys and yeah i am in flvs

OpenStudy (anonymous):

Hello DanJS

OpenStudy (danjs):

hi

OpenStudy (anonymous):

Do you think you can help me with another problem not this one but another or you don't know this stuff very well ?

OpenStudy (danjs):

sure, post it here and ill help you through it

OpenStudy (anonymous):

ok thanks hold on let me get it

OpenStudy (anonymous):

Solve the following system of equations: -2 + y = 1 -4x + y = -1 A.(3,1) B.(-1,3) C.(-1,-3) D.(1,3)

OpenStudy (anonymous):

Im ahead of you in the same class I can help you with anything that you need

OpenStudy (danjs):

If they give you multiple choice like that, you can simply test each solution without actually solving, if you want

OpenStudy (danjs):

for (x,y) = (3,1) , put into each equation and test if they are both true

OpenStudy (anonymous):

what lesson are you in and is your teacher ms.lamb

OpenStudy (anonymous):

No mine is Kelly Setta and thanks DanJS you really helped me thanks for your time

OpenStudy (anonymous):

And my lesson is

OpenStudy (anonymous):

Lesson 3: Solving Systems of Equations Algebraically

OpenStudy (danjs):

or if you would like to solve, there are a couple ways you can do that. Substitution: solve one equation for a single variable, and put tha into the next equation so you can solve an equation with a single variable. Combination: Multiply each equation through by a scalar number so that when you add the equations , one variable drops out, IE multiply the first equation by (-2) so that 4x +(-4x) = 0 and the x terms drop out

OpenStudy (danjs):

Or you can graph both of the lines, and the solution will be the intersection point (x,y)

OpenStudy (anonymous):

are you savannah ray

OpenStudy (anonymous):

My name ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

No why ?

OpenStudy (anonymous):

I thought it was because you can check who is in flvs currently so.........

OpenStudy (anonymous):

oh do you have this lesson your grade book also ?

OpenStudy (anonymous):

I think the graph option would be the easiest

OpenStudy (danjs):

What is FLVS?

OpenStudy (anonymous):

Florida Virtual School online school

OpenStudy (danjs):

Is that like ahigh school , or online college or...?

OpenStudy (anonymous):

High school

OpenStudy (danjs):

ah ic.. so you can do entire high school online these days?

OpenStudy (anonymous):

Yeah LOL

OpenStudy (danjs):

dang, how do they know it is You that is actually doing the stuff , rather than someone on your behalf? haha

OpenStudy (danjs):

Do you ever have to actually show up in person to anything? like a standardized test?

OpenStudy (anonymous):

They don't LOL but i do this a lot of times legit i just am stuck on this part

OpenStudy (anonymous):

and no

OpenStudy (danjs):

wow. ok. here.....

OpenStudy (danjs):

-2x + y = 1 -4x + y = -1 Multiply First equation through by (-2) 4x - 2y = -2 -4x + y = -1

OpenStudy (danjs):

Now add them together 4x + (-4x) + (-2y) + y = -2 + (-1) 0x - y = -3 y = +3

OpenStudy (danjs):

If y = 3, then from equation 1, -2(x) + 3 = 1 -2x = -2 x = +1

OpenStudy (danjs):

The solution is (x,y) = (1, 3)

OpenStudy (anonymous):

i see now they explain it so complicating you just explain it so simple thank you so very much

OpenStudy (danjs):

You can test that by putting in x=1 and y=3 into both equations, and see if they both are true

OpenStudy (danjs):

your welcome

OpenStudy (triciaal):

and the 4th method used to solve simultaneous equation is matrices. The combination method above is actually elimination and sometimes called elimination by combination. you are choosing the factors necessary to eliminate one variable when you combine the equation then solve for the one variable.